OpenStudy (anonymous):

Can someone help me factor this expression? ac + 2ad + 2bc + 4bd

OpenStudy (mertsj):

Group the first two terms and the last two terms and factor out the common factor. That is the first step. Can you do that?

OpenStudy (nenadmatematika):

=a(c+2d)+2b(c+2d)=(c+2d)(a+2b)

OpenStudy (anonymous):

ac + 2ad + 2bc + 4bd = (ac + 2ad) + (2bc + 4bd) = a(c + 2d) + 2b(c + 2d) = (a + 2b)(c + 2d)